OpenStudy (anonymous):

What inverse variation equation for the relationship. y varies inversely with x and y=5 when x=2. a. y=1/2x b. y=2x c. y=x+2 d. y=10/x

jhonyy9 (jhonyy9):

do you know what mean invers variation ?

jhonyy9 (jhonyy9):

or first can you tel me please what mean direct variation ?

jhonyy9 (jhonyy9):

ok looke plaese y=x mean that y variate directly with x - yes ? so how you think than what will be the invers variation of y to x ?

jhonyy9 (jhonyy9):

so invers variation mean that when y increasing , x will decreasing ok ?

jhonyy9 (jhonyy9):

do you understand how i have wrote there above ?

OpenStudy (jdoe0001):

hmmm I notice only one person saying anything at all

jhonyy9 (jhonyy9):

this mean that y= 1/x ok jdoe than please correct me

OpenStudy (jdoe0001):

either the poster fainted or wait to eat dinner or it's just waiting for an answer to transcribe over

OpenStudy (jdoe0001):

went to eat rather either way.... the poster is suspiciously quiet

jhonyy9 (jhonyy9):

@glamourglory do you understand anything from these ?

OpenStudy (anonymous):

To understand this question fully, it's pretty important to get what direct variation is. This is a direct variation function: y = 3x. (3 is, of course, a constant.) If x increases, y increases threefold. When both variables in the function increase, the variation can be said to be direct. So what makes an inverse variation function? Well, logic follows that if: direct variation = both variables increasing with change then: inverse variation = one variable increasing, while the other variable decreases. To represent the inverse nature of this function, we write the function: y = 3 / x. You can plug in values to see how this is an inverse function: If y = 1, x = 3 If y = 2, x = 1.5 If y = 3, x = 1 So we see the inverse nature of the function! Now, remembering the format of the function, y = 3 /x, we can address your issue. D pops right out at me. Gotta be the answer. Rock on!